(y+2) (y-3)=6
3a^2=14a+5
x(a-3)-y(3-a)
22z-9z^2-z3
a^4+a^3-72a^2

Answer by stanbon(75887)

y+2) (y-3)=6
y^2-y-6=6
y^2-y-12=0
(y-4)(y+3)=0
y=4 or y=-3
-----------------
3a^2=14a+5
3a^2-14a-5=0
3a^2-15a+a-5=0
3a(a-5)+(a-5)=0
(a-5)(3a+1)=0
a=5 or a=-1/3
---------------
x(a-3)-y(3-a)
=x(a-3)+y(a-3)
=(a-3)(x+y)
---------------

22z-9z^2-z^3
=-z(z^2+9z-22)
=-z(z+11)(z-2)
------------
a^4+a^3-72a^2
=a^2(a^2+a-72)
=a^2(a+9)(a-8)
Cheers,
Stan H.